OpenStudy (anonymous):

3x-7y=14 write in slope-intercept form

OpenStudy (neo92):

slope intercept form :- y = mx + c , m = slope , c = gradient just rearrange da gvn eqn 3x - 7y = 14 7y = 3x - 14 y = (3/7)x - 2

OpenStudy (neo92):

@pkd8 got it?

OpenStudy (anonymous):

got it ty Neo92